How many digits does a printer print if it prints whole numbers from 1 to 728, inclusive?

Answer by Theo(13342)

Should be:

1 to 9 = 9 - 1 + 1 = 8 + 1 = 9 * 1 digit = 9
10 to 99 = 99 - 10 + 1 = 89 + 1 = 90 * 2 digits = 180
100 to 728 = 728 - 100 + 1 = 628 + 1 = 629 * 3 digits = 1887

total = 9 + 180 + 1887 = 2076

As far as I can tell.